Valve configuration: Vacuum Normal is reached in really 4 days if we do not count overnight rest of roughing.
VA6 and VC2 are reconnected. I'm turning on the RGA next
All 4 ion pumps were vented with air and pumped down to ~ 1e-4 Torr
Ion pumps gate valve control cables are connected and their pumps are disconnected.
The black relay box was tested repeatedly and it stopped misbehaiving.
We were at atmosphere for 13 days. Chamber BS, ITMX, ITMY and ETMY were opened.
Al foil "cups" were placed on the back side OSEMs of PRM.
|